sir when we provw 2+2=5 then we write 16 in the form of ((2+2)2) but it is wrong because it is 8 plzzzzz tell me the right answer

sir


when we provw 2+2=5 then we write 16 in the form of ((2+2)2) but it is wrong because it is 8


plzzzzz tell me the right answer

Grade:11

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Baba

to prove 4=5

-4x5=-4x5

-20=-20

42 - 2x4x9/2 + (9/2)2 = 52 - 2x5x9/2 + (9/2)2

(4-9/2)2 = (5-9/2)2

4=5

All the best

AKASH GOYAL
AskiitiansExpert-IITD

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.